# A Survey on Class-Agnostic Counting: Advancements from Reference-Based to Open-World Text-Guided Approaches

### v3

@@ III-C Reference-less Class-agnostic Counting
The first reference-less class-agnostic counting method was introduced in [76]. In this paper, Ranjan et al. introduced the Repetitive Region Proposal Network Counter (RepRPN-Counter), which predicts separate density maps for each repetitive object category in a given query image. The model is based on FamNet [37] and includes an upstream module called the Repetitive Region Proposal Network (RepRPN), designed to automatically identify exemplars of repetitive objects in the image, which are then passed to FamNet for further processing. RepRPN takes inspiration from the Region Proposal Network (RPN) of the popular Faster R-CNN object detector [109]. RPN typically calculates proposal boxes and objectness scores, but RepRPN additionally outputs a so-called repetition score, indicating how frequently an object within the proposals appears in the image. These repetition scores are used to select exemplars from the repetitive object categories in the query image, with the highest-scoring proposals chosen as exemplars. However, the original RPN generated region proposals based on anchors – predefined bounding boxes of varying sizes and aspect ratios used as reference points for object detection – which did not cover the entire image, a critical requirement for computing the global repetition score. To address this, the authors incorporated Encoder Self-Attention layers [106], which can integrate information from similar feature vectors across the entire convolutional feature map of the image, enabling efficient computation of the global repetition score at any anchor location. Another challenge the authors faced was training RepRPN, particularly due to the missing labels in class-agnostic counting datasets such as FSC-147 [37]. Specifically, exemplar bounding boxes and dots are provided for only a single object category, but images may contain other unlabeled categories (see also Sec. IV). However, forcing RepRPN-Counter to predict zero counts for these unannotated objects could negatively impact its performance. To address this issue, the authors employed knowledge transfer from teacher networks. They used RepRPN, trained on the MSCOCO dataset [110], as a teacher network to generate target labels for the objectness and repetition scores of proposals that do not overlap with annotated objects in CAC datasets. Instead, to generate target density maps for the unannotated proposals, they leveraged a pre-trained FamNet, capable of predicting density maps for novel object classes based on a single exemplar.
@@
[77] went a step further by proposing RCC - Reference-less class-agnostic counting. Unlike previous works, it did not use point-level supervision – besides not using any exemplars as in the reference-less setting – thus proposing a weakly-supervised methodology. Specifically, their proposed network directly regresses to a single scalar count prediction for the entire input image, thus not requiring ground-truth density maps for supervised learning. To create a general and informative feature space without labels, the authors implemented a self-supervised knowledge distillation technique based on [111]. They encouraged the learning of meaningful image representations by fostering agreement between a fixed teacher network, which processes large global image crops, and a student network, which analyzes smaller local image crops. This alignment is achieved by minimizing the cross-entropy between the probability distributions produced by both networks.
@@
Another milestone concerning reference-less class-agnostic counting is represented by Counting TRansformer - CounTR [78] (pronounced "counter"). In this work, Liu et al. built on the idea that self-similarity serves as a strong prior in visual object counting. They introduced a transformer-based architecture where self-similarity can be explicitly captured by the inherent attention mechanisms, both between input image patches and with the exemplars (if any). Specifically, this architecture features two visual encoders: (i) a Video Transformer (ViT) [112], which maps input image features into a high-dimensional feature space, and (ii) a lightweight CNN, responsible for extracting visual features from the exemplars (if available). The subsequent feature interaction module is responsible for merging information from both encoders. It consists of a series of standard transformer decoder layers, where the image features serve as the Query, and two separate linear projections of the exemplar features, or a learnable special token if no exemplars are present, act as the Key and Value. This architecture aligns well with the self-similarity prior in counting tasks: self-attention captures similarities between image regions, while the cross-attention between Query and Value compares these regions with the exemplars or learns to bypass the CNN branch when using the learnable token, i.e., in the reference-less setting where no exemplars are provided. Additionally, the authors proposed a two-stage training approach, where the transformer-based image encoder is first pre-trained using self-supervision through masked image modeling [113], followed by supervised fine-tuning for the counting task. Finally, another notable contribution tackled the long-tailed issue in existing counting datasets, where only a few images contain a large number of instances. To handle this, they introduced a scalable pipeline for synthesizing training images with large object counts by cropping random-sized square regions from existing images and arranging them together in a collage.
@@
Also the authors of [79] introduced a technique that does not use point-level supervision such as [77], named GCNet - Generalized Counting Network. However, differently from [77], GCNet uses a pseudo-Siamese structure to capture pseudo-exemplar cues from the resized raw image. Specifically, the main branch of the network extracts primary 2D feature maps from an input image $I_{i}$. In contrast, the secondary branch is in charge of generating exemplar-focused outputs by exploiting a methodology inspired by vision transformers [112]: (i) a sequence of overlapping patches is extracted using the unfold operation over the 2D feature maps computed by the main branch, (ii) pseudo-exemplar patches are computed using a per-pixel average operation along the patch dimension, ensuring to capture the common attributes across local regions, (iii) pseudo-exemplar patches are divided into 16 sub-patches of size $2\times 2$, and (iv) each sub-patch is flattened into a 1D token and then transformed through linear projection. The output of the main branch is then further processed through an anisotropic encoder to capture discriminative features along three directions in feature space – horizontal, vertical, and basis/channel directions. Finally, these two outputs are fed into a dual-attention condenser. The same authors proposed also a reference-based version of GCNet, which attaches an additional ResNet-50 backbone to the front-end of the original GCNet for processing labeled exemplar patches; the exemplar features are then fed into the same back-end structure of GCNet to predict an auxiliary single count.
@@ III-D Open-world Text-guided Object Counting
Xu et al. [80] were the first to propose a open-world text-guided approach. They highlighted that the earlier reference-based paradigm is not well-suited for autonomous systems, as it relies on human intervention – requiring users to provide exemplar inputs for new categories during inference. Although reference-less class-agnostic counting eliminates the need for bounding box exemplars, it still has limitations in this context. Specifically, it does not allow for user interaction and only estimates the count of objects belonging to the dominant class(es) without identifying the class type. Consequently, its use in autonomous systems is limited to scenarios where the target object class aligns with the majority class(es) present among the objects. Thus, [80] introduces a more versatile paradigm called zero-shot counting (ZSC) where users can specify textual descriptions of the object classes they want to count, representing a sort of trade-off between previous class-agnostic counting paradigms – humans are still in the loop but can specify classes with less effort and in a more natural way. In more detail, their proposal is grounded on an effective strategy that selects meaningful patches in the input image containing the target object specified by the class name to serve as exemplars. To achieve this, they introduced a two-stage technique that (i) builds a class prototype in a pre-trained embedding space based on the given class name and selects patches whose embeddings are the $k$-nearest neighbors to that prototype, and (ii) filters out poor exemplars, retaining only the most relevant ones. Concerning the first step, the authors trained a conditional variational autoencoder (VAE) based on the popular Contrastive Language–Image Pretraining (CLIP) model [108] to generate visual exemplar prototypes conditioned with their semantic embedding encoded in the provided category name. The second stage relies instead on the observation that feature maps from meaningful exemplars show subtle differences compared to those generated from poor exemplars. Specifically, meaningful exemplars tend to display repetitive patterns centered around the object areas, while poor exemplars often exhibit irregular patterns throughout the image. To assess the quality of the generated feature maps – and thus the selected input exemplar patches – the authors employed a pre-trained reference-based object counter similar to FamNet [37], training it to predict the counting error associated with the input exemplar. Once this error predictor was trained, the authors utilized it to filter the selected patches, retaining only those with the smallest predicted errors, and to compute the density maps and associated instance counts in the input images for these exemplars.
@@
Concurrently, Amini-Naieni et al. [81] proposed an alternative solution that, unlike the two-stage pipeline in [80], introduced a single-stage approach called CounTX (pronounced "Count-text"). CounTX can be trained end-to-end and accepts more detailed specifications of target objects for counting, rather than relying solely on class names. Specifically, it extends the reference-less CounTR architecture [78] to the open-world text-guided setting. Like CounTR, it features a transformer-based image encoder, but it utilizes the CLIP vision transformer B-16 [108], pre-trained in a contrastive manner with a transformer-based text encoder using image-text pairs from LAION-2B [114]. In more detail, the image encoder encodes each input image into a spatial map of 512-dimensional feature vectors, while the text encoder processes the class description into a single 512-dimensional feature vector. To combine these features, the authors employed a transformer-based feature interaction module, similar to CounTR, but instead of matching image patches to visual exemplars, in CounTX, it is exploited to compute the similarities between image patches and class descriptions. Finally, the output of the feature interaction module is reshaped to a spatial feature map and fed into a progressive four-block convolution and upsampling operation to produce density maps.
@@
In [83], the authors incorporated three modules to efficiently fine-tune CLIP for the counting task and to exploit intermediate features across different encoding layers of CLIP in the decoding stage. In particular: (i) a semantic-conditioned prompt-tuning module is responsible for fine-tuning CLIP efficiently, using conditioning via semantic embeddings to generate patch embeddings that highlight the region of interest, rather than relying on simple learnable prompts; (ii) a learnable affine transformation module adjusts the semantic-patch similarity map produced by the previous module to better suit the counting task; (iii) a segment-aware skip connections module is utilized to capture intermediate features from various encoding layers of CLIP, enhancing the decoder’s generalization ability and providing rich contextual information.
@@
[84] addressed a key limitation in object counting, where the task is typically confined to the class level and fails to account for fine-grained details within the class. In practical applications, counting often requires contextual or referring human input to target specific objects, such as monitoring traffic flow in different directions or distinguishing between pedestrians and vehicles that are stationary or moving at various points of a junction. To tackle this, the authors introduced a novel task called Referring Expression Counting (REC), which focuses on counting objects with distinct attributes within the same class. To evaluate the REC task, they developed a new dataset named REC-8K, comprising 8011 images and 17,122 referring expressions. The methodology leverages the open-set detector GroundingDino [116], adapted with specific modifications such as replacing bounding boxes with box centers and utilizing the CLS token as a representation of the global semantics of the referring expression (rather than individual text tokens). Additionally, the authors introduced two modules to enhance REC performance: (i) a global-local feature fusion that integrates global image features to infer relational attributes, improving the model’s understanding of the context; (ii) a contrastive learning module that is designed to learn discriminative features associated with attributes by processing multiple referring expressions for different attributes of same-class objects. Besides the REC task, the authors also evaluated their method of class-agnostic counting to show that their approach is generalizable to prior tasks.
@@
In [85], the authors focused on contrasting the challenge of identifying high-quality exemplars effectively. Specifically, they presented the Visual Association-based Zero-shot Object Counting (VA-Count) framework, designed to establish a robust connection between object categories and their visual representations, enabling adaptability across various classes. VA-Count incorporates two key components: the Exemplar Enhancement Module (EEM) and the Noise Suppression Module (NSM), which respectively refine exemplar identification and mitigate adverse effects. The EEM leverages GroundingDino [116], as in [84]. However, here, the authors also introduced a binary filter to refine the candidate exemplars computed by GroundingDino, selecting the ones containing precisely only one object. However, despite the accurate selection of single-object exemplars, errors can arise when exemplars do not belong to the target category, introducing uncertainty in the learning process. The NSM acts as a safeguard by identifying and filtering out such negative exemplars. While the EEM focuses on selecting high-quality exemplars to establish strong visual associations, the NSM uses samples from irrelevant categories to refine these associations further. By employing contrastive learning, the NSM differentiates between positive exemplars (related to the target category) and negative exemplars, reducing the impact of misalignment on the counting task.
@@
Some reference-based techniques have been adapted for the open-world text-guided setting. For example, the authors of DAVE [72] introduced minor adjustments to their architecture, particularly in the cluster selection process during the verification stage. In this approach, the text prompt embedding is derived using the Contrastive Language–Image Pretraining (CLIP) model [108] and compared to the CLIP embeddings of the identified clusters. Cluster embeddings are computed by masking out image regions outside the bounding boxes associated with each cluster and then obtaining the CLIP embedding for the remaining regions. Cosine distances are calculated between the text embedding and each cluster embedding, and clusters with less than 85% of the highest prompt-to-cluster similarity score are classified as outliers. Similarly, the authors of TFPOC [73] enabled textual inputs to define object classes. Their method begins with a coarse similarity map generated using CLIP-Surgery [117], an enhanced variant of CLIP. From this map, reference objects are identified through a series of steps: the similarity map is binarized, the largest connected component (likely containing the target objects) is isolated, and bounding boxes are created around sub-regions within this component for further refinement. These bounding boxes are then used as prompts for Segment Anything Model [107], producing a high-quality similarity map. In another approach, the authors of [74] employed CLIP to generate a prompt mask inspired by MaskCLIP [118]. Cosine distance was utilized to compute the matching score between predefined text features and the corresponding visual features. In PseCo [75], similar to the few-shot setting, the authors used the CLIP text embeddings as classification weights when only known class names.
@@ IV Datasets
There is a lack of publicly available data for the CAC task. Indeed, most existing counting datasets are designed for specific object categories. Notable examples include UCF-QNRF [119], ShanghaiTech [86], and WorldExpo’10 [4] for crowd counting; TRANCOS [120], NDISPark [42, 121], and CARPK [44] for vehicle counting; and VGG Cell [30] for cell estimation. Meanwhile, datasets that include multiple object categories, such as the MSCOCO dataset [110], are unsuitable for counting tasks because they were designed for object detection, typically featuring images with only a small number of object instances. In this section, we outline the limited CAC benchmarks available in the literature. Among them, FSC-147 [37] is considered the gold standard. Additionally, although originally intended for vehicle counting, CARPK [44] is widely used to evaluate generalization capabilities. Fig. 5 illustrates some samples from these two latter benchmarks.
@@
The FSC-147 dataset comprises 6,135 images spanning 147 object categories, including plants, animals, vehicles, and food. These images were sourced from Flickr, Google, and Bing using open-source image scrapers and then manually inspected and filtered. Annotations were manually created by marking dots over the approximate centroids of each object instance – as usual for the object counting task [1]. Occluded instances were annotated only if at least 10% of the object was visible. Additionally, the dataset includes the coordinates of three bounding boxes per image, representing exemplar instances. Each image contains exemplars belonging to a single category, whose natural language name is stored in a text file. If multiple categories were present in an image, one was chosen arbitrarily. Another characteristic of FSC-147 is the significant variation in object counts per image, ranging from 7 to 3,731, with an average of 56 objects per image. The dataset is divided into three splits: 89 object categories are allocated to the training set, while 29 categories are assigned to the validation and testing sets. These splits consist of 3,659, 1,286, and 1,190 images, respectively.
@@
A revised version of FSC-147, called FSC-133, was introduced in [77]. The authors identified and corrected several errors in FSC-147. Specifically, they found that (i) 11 images from the training set also appeared in the validation or testing sets, (ii) some images were effectively duplicates, exhibiting pixel-wise differences of 0 or near 0 when compared at a $224\times 224$ resolution, and (iii) certain duplicate images had inconsistent ground truth counts, with discrepancies of up to 25%. The data split overlaps stemmed from misclassification due to ambiguous or hierarchical categories in FSC-147, such as ‘kidney beans’ and ‘red beans’ or ‘bread rolls’ and ‘baguette rolls.’ In contrast, count discrepancies were observed in images with challenging object occlusions. To address these issues, FSC-133 merged very similar categories and retained only the images with the most accurate counts. The revised dataset consists of 5,898 images across 133 categories, with splits containing 3,877 images for training, 954 for validation, and 1,067 for testing. However, despite the improvements introduced with FSC-133, subsequent works have largely continued to use FSC-147, which remains the standard for this task.
FSC-147-D
Another enhanced version of FSC-147, called FSC-147-D, was introduced by [81] to better support the open-world text-guided object counting paradigm. In this version, Amini-Naieni et al. replaced the simpler object descriptions with more fine-grained and structured natural language sentences for each image, specifying the objects to be counted with greater precision. For example, the original FSC-147 class name for the image with identifier 3696 was updated from "candy pieces" to "the pink desserts garnished with candy pieces".
@@
Moreover, two datasets have addressed a key limitation of FSC-147 – its limited number of images featuring multiple object categories – by including images with multiple categories: OmniCount-191 [122] and MCAC [123]. However, OmniCount-191 is not publicly available, and MCAC has significant drawbacks: it lacks appropriate annotations for open-world text-guided approaches and is synthetic, failing to represent realistic scenarios.
@@
Finally, several existing papers have also evaluated their performance on the CARPK dataset [44], a collection of drone-view images of parking lots specifically designed for vehicle counting. This benchmark was first utilized for class-agnostic counting by the work that introduced FSC-147 and has since been used by subsequent works, primarily to test cross-dataset generalization. The dataset comprises 1,448 images, divided into a training set of 989 images and a test set of 459 images, with approximately 90,000 annotated vehicles in total. The annotations feature bounding boxes, and density maps are generated using dots placed at the centers of these bounding boxes. Additionally, some of the bounding boxes are utilized to identify exemplars.
Figure 5: Dataset samples. We report some examples of the CAC gold standard benchmarks. a) The FSC-147 dataset [37] is the standard for class-agnostic counting. It includes more than 6,000 images belonging to 147 object classes. We show some sample images together with the provided labels – bounding boxes localizing the exemplars and density maps. b) The CARPK dataset [44] is a set of drone-view images tailored for vehicle counting that is often exploited for assessing the generalization capabilities of CAC models. We show some sample images together with bounding boxes marking the exemplars.
V Results and Discussion
In this section, we present the results of the evaluated methods, following the same taxonomy of previous sections – reference-based, reference-less, and open-world text-guided approaches. First, we describe the experimental setting and then provide a critical analysis of the outcomes across various points of discussion.
V-A Experimental Setting
Some techniques have been adapted to carry out more than one of these tasks (see also Sec. III); in these cases, we report the results for all of them. We considered the two most commonly used benchmarks: FSC-147 [37] and CARPK [44] (see Table II and Table V-A). For FSC-147, in line with existing literature, we report results for both the validation and test subsets. Differently, the CARPK dataset is used to evaluate the models’ generalization capabilities in counting instances of a specific object class – vehicles. Previous works have explored two different settings: i) training models on FSC-147 without using any data from CARPK or the car category in FSC-147, and ii) fine-tuning models using training data from CARPK, where a set of 12 exemplars is randomly sampled from the training set and used as exemplars across all training and test images.
TABLE II: Results on the FSC-147 dataset [37]. We report the performance in terms of MAE and RMSE reached by the methodologies discussed in this survey over the gold standard CAC benchmark – FSC-147. The best results are in bold.

### v4

@@ III-C Reference-less Class-agnostic Counting
The first reference-less class-agnostic counting method was introduced in [79]. In this paper, Ranjan et al. introduced the Repetitive Region Proposal Network Counter (RepRPN-Counter), which predicts separate density maps for each repetitive object category in a given query image. The model is based on FamNet [32] and includes an upstream module called the Repetitive Region Proposal Network (RepRPN), designed to automatically identify exemplars of repetitive objects in the image, which are then passed to FamNet for further processing. RepRPN takes inspiration from the Region Proposal Network (RPN) of the popular Faster R-CNN object detector [108]. RPN typically calculates proposal boxes and objectness scores, but RepRPN additionally outputs a so-called repetition score, indicating how frequently an object within the proposals appears in the image. These repetition scores are used to select exemplars from the repetitive object categories in the query image, with the highest-scoring proposals chosen as exemplars. A key challenge in training RepRPN was the lack of labels in class-agnostic counting datasets like FSC-147 [32], where annotations are provided for only one object category, despite the presence of other unannotated objects (see also Sec. IV). Penalizing predictions for these unlabeled objects could harm performance. To overcome this, the authors used knowledge transfer: RepRPN trained on MSCOCO [109] acted as a teacher to assign objectness and repetition scores to proposals not overlapping with annotated objects, while a pre-trained FamNet generated density maps for these proposals using single exemplars.
@@
[80] went a step further by proposing RCC - Reference-less class-agnostic counting. Unlike previous works, it did not use point-level supervision—besides not using any exemplars as in the reference-less setting—thus proposing a weakly-supervised methodology. Specifically, their proposed network directly regresses to a single scalar count prediction for the entire input image, thus not requiring ground-truth density maps for supervised learning. To create a general and informative feature space without labels, the authors implemented a self-supervised knowledge distillation technique based on [110]. They encouraged the learning of meaningful image representations by fostering agreement between a fixed teacher network, which processes large global image crops, and a student network, which analyzes smaller local image crops. This alignment is achieved by minimizing the cross-entropy between the probability distributions produced by both networks.
@@
Another milestone concerning reference-less class-agnostic counting is represented by Counting TRansformer - CounTR [81] (pronounced "counter"). In this work, Liu et al. built on the idea that self-similarity serves as a strong prior in visual object counting. They introduced a transformer-based architecture where self-similarity can be explicitly captured by the inherent attention mechanisms, both between input image patches and with the exemplars (if any). Specifically, this architecture features two visual encoders: (i) a Video Transformer (ViT) [111], which maps input image features into a high-dimensional feature space, and (ii) a lightweight CNN, responsible for extracting visual features from the exemplars (if available). The subsequent feature interaction module is responsible for merging information from both encoders. It consists of a series of standard transformer decoder layers, where the image features serve as the Query, and two separate linear projections of the exemplar features, or a learnable special token if no exemplars are present, act as the Key and Value. This architecture aligns well with the self-similarity prior in counting tasks: self-attention captures similarities between image regions, while the cross-attention between Query and Value compares these regions with the exemplars or learns to bypass the CNN branch when using the learnable token, i.e., in the reference-less setting where no exemplars are provided. Additionally, the authors proposed a two-stage training approach, where the transformer-based image encoder is first pre-trained using self-supervision through masked image modeling [112], followed by supervised fine-tuning for the counting task.
@@
Also the authors of [82] introduced a technique that does not use point-level supervision such as [80], named GCNet - Generalized Counting Network. However, differently from [80], GCNet uses a pseudo-Siamese structure to capture pseudo-exemplar cues from the resized raw image. Specifically, the main branch of the network extracts primary 2D feature maps from an input image $I_{i}$. In contrast, the secondary branch is in charge of generating exemplar-focused outputs by exploiting a methodology inspired by vision transformers [111]. The output of the main branch is then further processed through an anisotropic encoder to capture discriminative features along three directions in feature space—horizontal, vertical, and basis/channel directions. Finally, these two outputs are fed into a dual-attention condenser. The same authors also proposed a reference-based version of GCNet, which attaches an additional ResNet-50 backbone to the front-end of the original GCNet for processing labeled exemplar patches; the exemplar features are then fed into the same back-end structure of GCNet to predict an auxiliary single count.
@@ III-D Open-world Text-guided Object Counting
Xu et al. [83] were the first to propose a open-world text-guided approach, introducing a more versatile paradigm called zero-shot counting (ZSC) where users can specify textual descriptions of the object classes they want to count, representing a sort of trade-off between previous class-agnostic counting paradigms—humans are still in the loop but can specify classes with less effort and in a more natural way. In more detail, their proposal is grounded on an effective strategy for selecting meaningful patches from the input image that contain the target object specified by a class name, to be used as exemplars. To this end, they construct a class prototype in a pre-trained embedding space using the provided class name, and then select patches whose embeddings are among the $k$-nearest neighbors to that prototype. This is implemented using a conditional variational autoencoder (VAE) built on the widely adopted Contrastive Language-Image Pretraining (CLIP) model [107], which generates visual exemplar prototypes conditioned on the semantic embedding of the given category name.
@@
Concurrently, Amini-Naieni et al. [84] proposed an alternative solution called CounTX (pronounced "Count-text"). CounTX can be trained end-to-end and accepts more detailed specifications of target objects for counting, rather than relying solely on class names. To support this, the authors introduced an enhanced version of the FSC-147 dataset [32], named FSC-147-D, where object classes are described using more fine-grained and structured natural language sentences (See also Sec. IV for more details). Specifically, CounTX extends the reference-less CounTR architecture [81] to the open-world text-guided setting. Like CounTR, it features a transformer-based image encoder, but it utilizes the CLIP vision transformer B-16 [107], pre-trained in a contrastive manner with a transformer-based text encoder using image-text pairs from LAION-2B [113]. In more detail, the image encoder encodes each input image into a spatial map of 512-dimensional feature vectors, while the text encoder processes the class description into a single 512-dimensional feature vector. To combine these features, the authors employed a transformer-based feature interaction module, similar to CounTR, but instead of matching image patches to visual exemplars, in CounTX, it is exploited to compute the similarities between image patches and class descriptions. Finally, the output of the feature interaction module is reshaped to a spatial feature map and fed into a progressive four-block convolution and upsampling operation to produce density maps.
@@
In [86], the authors incorporated three modules to efficiently fine-tune CLIP for the counting task and to exploit intermediate features across different encoding layers of CLIP in the decoding stage. In particular: (i) a semantic-conditioned prompt-tuning module is responsible for fine-tuning CLIP efficiently, using conditioning via semantic embeddings to generate patch embeddings that highlight the region of interest, rather than relying on simple learnable prompts; (ii) a learnable affine transformation module adjusts the semantic-patch similarity map produced by the previous module to better suit the counting task; (iii) a segment-aware skip connections module is utilized to capture intermediate features from various encoding layers of CLIP, enhancing the generalization ability of the decoder and providing rich contextual information.
@@
[87] addressed a key limitation in object counting, which is often restricted to class-level granularity and overlooks fine-grained distinctions. However, in real-world scenarios, counting typically requires contextual or referential input—for example, distinguishing traffic flow directions or separating stationary from moving pedestrians and vehicles at a junction. To tackle this, the authors introduced a novel task called Referring Expression Counting (REC), which focuses on counting objects with distinct attributes within the same class. To evaluate the REC task, they developed a new dataset named REC-8K, comprising 8011 images and 17,122 referring expressions (See also Sec. IV for more details). The methodology leverages the open-set detector GroundingDino [115], adapted with specific modifications such as replacing bounding boxes with box centers and utilizing the CLS token as a representation of the global semantics of the referring expression (rather than individual text tokens). In addition to REC, they also evaluated their method on class-agnostic counting, demonstrating its generalizability to prior tasks.
@@
In [88], the authors addressed the challenge of identifying high-quality exemplars for zero-shot counting, introducing the Visual Association-based Zero-shot Object Counting (VA-Count) framework. VA-Count includes two main components: the Exemplar Enhancement Module (EEM) and the Noise Suppression Module (NSM), which respectively refine exemplar selection and mitigate the impact of incorrect samples. The EEM leverages GroundingDino [115], as in [87], but adds a binary filter to select candidate exemplars containing exactly one object. Despite this, errors may occur when selected exemplars do not belong to the target category. The NSM addresses this by identifying and filtering out such negative exemplars. Using contrastive learning, it distinguishes between relevant and irrelevant samples, enhancing the robustness of the counting process.
Figure 5: Dataset samples. We show some samples from the CAC gold standard benchmarks, together with bounding boxes marking the exemplars. (a) The FSC-147 dataset [32] includes more than 6,000 images belonging to 147 object classes. We present sample images along with the provided labels, which include density maps and three bounding boxes localizing the exemplars. (b) The CARPK dataset [43], a collection of drone-view images tailored for vehicle counting, which is often used to assess cross-dataset generalization capabilities on the specific vehicle category. We show sample images together with 12 bounding boxes marking the exemplars (as set by [32]).
@@
Some reference-based techniques have been adapted for the open-world text-guided setting. For example, the authors of DAVE [74] introduced minor adjustments to their architecture, particularly in the cluster selection process during the verification stage. In this approach, the text prompt embedding is derived using the CLIP model [107] and compared to the CLIP embeddings of the identified clusters. Cluster embeddings are computed by masking out image regions outside the bounding boxes associated with each cluster and then obtaining the CLIP embedding for the remaining regions. Cosine distances are calculated between the text embedding and each cluster embedding, and clusters with less than 85% of the highest prompt-to-cluster similarity score are classified as outliers. Similarly, the authors of TFPOC [75] enabled textual inputs to define object classes. Their method begins with a coarse similarity map generated using CLIP-Surgery [116], an enhanced variant of CLIP. From this map, reference objects are identified through a series of steps: the similarity map is binarized, the largest connected component (likely containing the target objects) is isolated, and bounding boxes are created around sub-regions within this component for further refinement. These bounding boxes are then used as prompts for Segment Anything Model [106], producing a high-quality similarity map. In another approach, the authors of [76] employed CLIP to generate a prompt mask inspired by MaskCLIP [117]. Cosine distance was utilized to compute the matching score between predefined text features and the corresponding visual features. In PseCo [77], similar to the few-shot setting, the authors used the CLIP text embeddings as classification weights when only known class names. Finally, CountDiff [78] can be straightforwardly switched to perform open-world text-guided CAC by obtaining the object-specific embedding leveraging the pre-trained text encoder in the text-to-image diffusion model
@@ IV Datasets
There is a lack of publicly available data for the CAC task. Indeed, most existing counting datasets are designed for specific object categories. Notable examples include UCF-QNRF [118], ShanghaiTech [59], and WorldExpo’10 [4] for crowd counting; TRANCOS [119] and NDISPark [41, 120] for vehicle counting; Pest Sticky Traps [18, 121] for pest abundance estimation; and VGG Cell [31] for cell estimation. Meanwhile, datasets that include multiple object categories, such as the MSCOCO dataset [109], are unsuitable for counting tasks because they were designed for object detection, typically featuring images with only a small number of object instances. In this section, we outline the limited CAC benchmarks available in the literature. Among them, FSC-147 [32] is considered the gold standard. Additionally, although originally intended for vehicle counting, CARPK [43] is widely used to evaluate generalization capabilities. Fig. 5 illustrates some samples from these two latter benchmarks.
@@
The FSC-147 dataset comprises 6,135 images spanning 147 object categories, including plants, animals, vehicles, and food. Annotations were manually created by marking dots over the approximate centroids of each object instance—as usual for the object counting task [1]. Additionally, the dataset includes the coordinates of three bounding boxes per image, representing exemplar instances. Each image contains exemplars belonging to a single category, whose natural language name is stored in a text file. If multiple categories were present in an image, one was chosen arbitrarily. Another characteristic of FSC-147 is the significant variation in object counts per image, ranging from 7 to 3,731, with an average of 56 objects per image. The dataset is divided into three splits: 89 object categories are allocated to the training set, while 29 categories are assigned to the validation and testing sets. These splits consist of 3,659, 1,286, and 1,190 images, respectively.
@@
A revised version of FSC-147, called FSC-133, was introduced in [80]. The authors identified and corrected several errors in FSC-147. Specifically, they found that (i) 11 training images also appeared in validation or test sets, (ii) some images were near-duplicates with pixel-wise differences close to zero, and (iii) certain duplicates had inconsistent ground truth counts, with discrepancies up to 25%. To address these issues, FSC-133 merged very similar categories and retained only the images with the most accurate counts. The revised dataset consists of 5,898 images across 133 categories, with splits containing 3,877 images for training, 954 for validation, and 1,067 for testing. However, despite the improvements introduced with FSC-133, subsequent works have largely continued to use FSC-147, which remains the standard for this task.
FSC-147-D and REC-8K
Another enhanced version of FSC-147, called FSC-147-D, was introduced by [84] to better support the open-world text-guided object counting paradigm. In this version, Amini-Naieni et al. replaced the simpler object descriptions with more fine-grained and structured natural language sentences for each image, specifying the objects to be counted with greater precision. In a similar direction, REC-8K [87] is a collection of approximately 8,000 images drawn from existing datasets, including FSC-147, where the textual descriptions of object categories are enriched with attributes.
@@
Moreover, two datasets have addressed a key limitation of FSC-147—its limited number of images featuring multiple object categories—by including images with multiple categories: OmniCount-191 [122] and MCAC [123]. However, both present significant drawbacks. OmniCount-191 contains a low number of objects per image, and many images within the same category are visually similar—often being augmented versions of the same source—even across training and test splits. MCAC, on the other hand, is synthetic and lacks appropriate annotations for open-world text-guided approaches, limiting its applicability to realistic scenarios. Notably, still within the context of synthetic data, [124] and [125] proposed approaches relying on text-to-image latent diffusion models to produce counting data across diverse object categories.
@@
Finally, several existing papers have also evaluated their performance on the CARPK dataset [43], a collection of drone-view images of parking lots specifically designed for vehicle counting. This benchmark was first utilized for CAC by the work that introduced FSC-147 and has since been used by subsequent works, primarily to assess cross-dataset generalization on the specific vehicle category. The dataset comprises 1,448 images, divided into a training set of 989 images and a test set of 459 images, with approximately 90,000 annotated vehicles in total. In CAC, this dataset is used either without any annotations or with a set of 12 randomly selected exemplars.
TABLE II: Results on the FSC-147 dataset [32]. We report the performance in terms of MAE and RMSE reached by the methodologies discussed in this survey over the gold standard CAC benchmark—FSC-147. The best results are in bold.